OpenStudy (anonymous):

1+ i : write the complex number in polar form with argument theta between 0 an 2 pi.... I know the answer but i need help getting to it

OpenStudy (accessdenied):

What did you have for an answer? And have you tried anything in particular so far?

OpenStudy (accessdenied):

We want to find polar form, which is a combination of a 'radius'/magnitude and angle from the terminal axis (the positive real line). We currently know its rectangular equivalent is z = 1 + i .

OpenStudy (anonymous):

I found out how.. now im just stuck on another problem

OpenStudy (accessdenied):

Oh, alright. If you require assistance on it, feel free to post the problem here. :p

OpenStudy (anonymous):

haha okay i need help writing 2 sqrt 3 - 2i in polar form

OpenStudy (anonymous):

i have the argument down but i do not know how to square the 4 radical 3

OpenStudy (accessdenied):

As in: \( r = \left( 2 \sqrt{3} \right)^2 + 2^2\) ?

OpenStudy (anonymous):

yes!

OpenStudy (accessdenied):

You could use the property of exponents of distributing to each factor: \(\left(ab \right)^x = a^x b^x \) \( \left(2 \sqrt{3} \right)^2 = 2^2 \left(\sqrt{3}\right)^2 \) Does that help?

OpenStudy (anonymous):

oh wait 12

OpenStudy (accessdenied):

Yeah. The square and square root cancel. :)

OpenStudy (anonymous):

but the answer says it is \[4( \cos 11\pi/6 + i \sin 11\pi/6)\]

OpenStudy (accessdenied):

Oops, I made a small mistake above with that radius formula: \( r^{\color{red}2} = \left(2\sqrt{3}\right)^2 + 2^2 \)

OpenStudy (accessdenied):

r^2 = 12 + 4 = 16 so r = sqrt(16) or 4.

OpenStudy (anonymous):

oh wow okay!!

OpenStudy (anonymous):

thank you

OpenStudy (accessdenied):

Yep, you're welcome! :)